how to convert 50o 37’ 30’’ into radians ?

Grade:12

5 Answers

Abbas
8 years ago
Hi Gaurav , first of all you ned
Abbas
8 years ago
Hi Gaurav , first of all you need to divide 30” b y 60 to convert it into minutes (as 1’ = 60”) , now add the answer you got(i.e 0.5) to 37’ as they both are  minutes terms , after this again diviide this new minute term (i.e 37.5’ ) by 60 to convert it into degrees (as 10 = 60’), now again add the answer (i.e 0.625) to 500 as they both are in degrees . Now the answer should be 50.6250 , this is in degrees , finally to convert it into radians multiply this answer with 3.14 and divide by 180 (as 10 = 3.14/180 radians ) . The answer will come out to be 0.883125 radians . Note if you want answer with pi as it is then don’t multiply it with 3.14 just divide it by 180. So your answer with pi would be 0.28125*pi radians . I hope you got your answer..:)
 
 
 
Abbas
8 years ago
Sorry for the first comment it was by mistake...don’t disapprove it or this one..
 
 
 
Prajwal Kavad
8 years ago
just use simple concept that pie radians =180degree
saniakhan
6 years ago
USe the Concept of Pie Radians=180 degree
Abbas
 
 
 
 
Hi Gaurav , first of all you need to divide 30” b y 60 to convert it into minutes (as 1’ = 60”) , now add the answer you got(i.e 0.5) to 37’ as they both are  minutes terms , after this again diviide this new minute term (i.e 37.5’ ) by 60 to convert it into degrees (as 10 = 60’), now again add the answer (i.e 0.625) to 500 as they both are in degrees . Now the answer should be 50.6250 , this is in degrees , finally to convert it into radians multiply this answer with 3.14 and divide by 180 (as 10 = 3.14/180 radians ) . The answer will come out to be 0.883125 radians . Note if you want answer with pi as it is then don’t multiply it with 3.14 just divide it by 180. So your answer with pi would be 0.28125*pi radians . I hope you got your answer..:)
Abbas
 
 
 
 
Hi Gaurav , first of all you need to divide 30” b y 60 to convert it into minutes (as 1’ = 60”) , now add the answer you got(i.e 0.5) to 37’ as they both are  minutes terms , after this again diviide this new minute term (i.e 37.5’ ) by 60 to convert it into degrees (as 10 = 60’), now again add the answer (i.e 0.625) to 500 as they both are in degrees . Now the answer should be 50.6250 , this is in degrees , finally to convert it into radians multiply this answer with 3.14 and divide by 180 (as 10 = 3.14/180 radians ) . The answer will come out to be 0.883125 radians . Note if you want answer with pi as it is then don’t multiply it with 3.14 just divide it by 180. So your answer with pi would be 0.28125*pi radians . I hope you got your answer..:)
 
 
 
 
 
Hi Gaurav , first of all you need to divide 30” b y 60 to convert it into minutes (as 1’ = 60”) , now add the answer you got(i.e 0.5) to 37’ as they both are  minutes terms , after this again diviide this new minute term (i.e 37.5’ ) by 60 to convert it into degrees (as 10 = 60’), now again add the answer (i.e 0.625) to 500 as they both are in degrees . Now the answer should be 50.6250 , this is in degrees , finally to convert it into radians multiply this answer with 3.14 and divide by 180 (as 10 = 3.14/180 radians ) . The answer will come out to be 0.883125 radians . Note if you want answer with pi as it is then don’t multiply it with 3.14 just divide it by 180. So your answer with pi would be 0.28125*pi radians . I hope you got your answer..:)
 
 
 
Hi Gaurav , first of all you need to divide 30” b y 60 to convert it into minutes (as 1’ = 60”) , now add the answer you got(i.e 0.5) to 37’ as they both are  minutes terms , after this again diviide this new minute term (i.e 37.5’ ) by 60 to convert it into degrees (as 10 = 60’), now again add the answer (i.e 0.625) to 500 as they both are in degrees . Now the answer should be 50.6250 , this is in degrees , finally to convert it into radians multiply this answer with 3.14 and divide by 180 (as 10 = 3.14/180 radians ) . The answer will come out to be 0.883125 radians . Note if you want answer with pi as it is then don’t multiply it with 3.14 just divide it by 180. So your answer with pi would be 0.28125*pi radians . I hope you got your answer..:)